# Birth and Death of a Rose

### v1

@@ Abstract
We study the problem of generating temporal object intrinsics—temporally evolving sequences of object geometry, reflectance, and texture, such as a blooming rose—from pre-trained 2D foundation models. Unlike conventional 3D modeling and animation techniques that require extensive manual effort and expertise, we introduce a method that generates such assets with signals distilled from pre-trained 2D diffusion models. To ensure the temporal consistency of object intrinsics, we propose Neural Templates for temporal-state-guided distillation, derived automatically from image features from self-supervised learning. Our method can generate high-quality temporal object intrinsics for several natural phenomena and enable the sampling and controllable rendering of these dynamic objects from any viewpoint, under any environmental lighting conditions, at any time of their lifespan. Project website: https://chen-geng.com/rose4d
@@ 1 Introduction
To summarize, we propose a new task of generating temporal object intrinsics, in the form of temporally-evolving sequences of 3D shape, reflectance, and texture. We introduce a framework to distill 4D-consistent temporal object intrinsics from pre-trained 2D diffusion models. Central to this framework is a canonical Neural Template that anchors the distillation signals to specific temporal states. We test this framework on several different object categories. We quantitatively compare its performance with prior art on automatic 4D generation, suggesting an advantage of the proposed methods across different examples. A further ablation study shows that the core modules and techniques proposed in this framework are crucial for performance.
@@ 2 Related Work
Figure 4: Internal structure of Neural Template. Taking camera pose $\xi$ and timestamp $t$ as input, the neural template first render a task-specific coarse deformable mesh under the provided condition. It is further denoised with a consistency model and encoded to a feature map, which will be further used to control the diffusion models.
@@ 3.2 Method Motivation
In fact, after spatial downsampling and Principal Component Analysis (PCA), 2D image feature maps extracted using recent 2D foundation models [37] can effectively provide such semantic affinity information, which is similar to the projection of skeleton in traditional animation. As shown in Figure 3, these feature maps effectively encode the temporal state as well as the viewpoint information for a natural process. We therefore call those maps “Neural State Maps”. We empirically found that these maps can serve as conditional signals for ControlNet [74] to control the temporal state and viewpoint of generated images.
@@ 3.3 System Overview
We build a system that takes in the prompt of a natural process, such as “a rose blooming”, and generate temporal object intrinsics, as shown in Figure 2. The generation process is split into two stages. In the first stage, we sample a reference video from a video diffusion model [66] and reconstruct a coarse deformable mesh to represent the 4D temporal stages in the natural process. The reconstructed coarse deformable mesh can be used to construct the Neural Template as in Figure 4, which will be used in the second stage. We discuss the details of this process in Sec. 4.
@@ 4 Neural Template
4.1 Estimating Coarse Deformable Mesh
@@ 5.1 Hybrid 4D Representation
We now introduce a 4D representation for $\mathcal{G}$ which is both expressive and temporally consistent. We adopt a hybrid 4D representation with K-Planes [13] representing low-frequency information and with Neural Graphical Primitives (NGP) [36] representing high-frequency details.
@@
where $\psi_{t}$ is the NGP encoding for timestamp $t$, and $t_{i}$ and $t_{i+1}$ are two keyframes around the sampled timestamp.
The full 4D feature $f_{\text{4D}}$ is computed by concatenating $f_{\text{low}}$ and $f_{\text{high}}$. We further pass the 4D feature into two MLPs to obtain signed distance values and material parameters for neural rendering [58, 35] as explained next.
@@ 5.2 Physically-Based Rendering
We now describe the rendering procedure for the hybrid representation of $\mathcal{G}$ from Sec. 5.1. Following common conventions, we use the Disney Principled PBR material model [7]. Specifically, the material parameters consist of diffuse color $k_{d}$, the visibility map $V(\mathbf{x})$, and $k_{\text{orm}}=(o,r,m)$ with roughness $r$, metallic parameter $m$, and $o$ left unused. The specular color $k_{s}$ is computed with $k_{s}=(1-m)\cdot 0.04+m\cdot k_{d}$. The outgoing radiance $L(\omega_{o})$ in direction $\omega_{o}$ can be represented with the rendering equation [22]:
@@
where $\mathbf{n}$ is the surface normal, $L_{i}(\mathbf{x},\omega_{i})$ is the incoming radiance from direction $\omega_{i}$, $f(\omega_{i},\omega_{o})$ is the BSDF computed from the material parameters. More details can be found in the supplementary material.
@@ 6.2 Comparison with Prior Works
Figure 8: Ablation studies on estimating the coarse dynamic mesh. The ARAP and the flow loss are both crucial for creating the coarse deformable mesh from the reference video.
@@ 6.3 Ablations and Validations
Learning Coarse Deformable Mesh. We study alternative approaches to obtain the coarse reference mesh in Neural Template. The results are in Figure 8. ARAP regularizes the learning of the deformation field to make it physically plausible. The flow loss helps to give a clearer gradient to the optimization of deformation fields. If we reconstruct the 3D shape frame-by-frame, the model will fail to give a clean back view due to the limitations of existing 3D reconstruction models.
@@ Acknowledgments.
We thank Haian Jin, Zizhang Li, and Ruocheng Wang for insightful discussions and feedback on the paper. This work is in part supported by NSF RI #2211258, #2338203, ONR MURI N00014-22-1-2740, and Samsung. YZ is in part supported by the Stanford Interdisciplinary Graduate Fellowship.

### v2

@@ Abstract
We study the problem of generating temporal object intrinsics—temporally evolving sequences of object geometry, reflectance, and texture, such as a blooming rose—from pre-trained 2D foundation models. Unlike conventional 3D modeling and animation techniques that require extensive manual effort and expertise, we introduce a method that generates such assets with signals distilled from pre-trained 2D diffusion models. To ensure the temporal consistency of object intrinsics, we propose Neural Templates for temporal-state-guided distillation, derived automatically from image features from self-supervised learning. Our method can generate high-quality temporal object intrinsics for several natural phenomena and enable the sampling and controllable rendering of these dynamic objects from any viewpoint, under any environmental lighting conditions, at any time of their lifespan. Project website: https://chen-geng.com/rose4d.
@@ 1 Introduction
To summarize, we propose a new task of generating temporal object intrinsics, in the form of temporally-evolving sequences of 3D shape, reflectance, and texture. We introduce a framework to distill 4D-consistent temporal object intrinsics from pre-trained 2D diffusion models. Central to this framework is a canonical Neural Template that anchors the distillation signals to specific temporal states. We test this framework on several different object categories. We quantitatively compare its performance with prior state of the art on automatic 4D generation, suggesting advantages of the proposed methods across different examples. A further ablation study shows that the core modules and techniques proposed in this framework are crucial for performance.
@@ 2 Related Work
Figure 4: Internal structure of Neural Template. Taking camera pose $\xi$ and timestamp $t$ as input, the neural template first render a task-specific coarse dynamic mesh under the provided condition. It is further denoised with a consistency model and encoded to a feature map, which will be further used to control the diffusion models.
@@ 3.2 Method Motivation
In fact, after spatial downsampling and Principal Component Analysis (PCA), 2D image feature maps extracted using DINOv2 [39] can effectively provide such semantic affinity information, which is similar to the projection of skeleton in traditional animation. As shown in Figure 3, these feature maps effectively encode the temporal state as well as the viewpoint information for a natural process. We therefore call those maps “Neural State Maps”. We empirically found that these maps can serve as conditional signals for ControlNet [77] to control the temporal state and viewpoint of generated images.
@@ 3.3 System Overview
We build a system that takes in the prompt of a natural process, such as “a rose blooming”, and generate temporal object intrinsics, as shown in Figure 2. The generation process is split into two stages. In the first stage, we construct a coarse dynamic mesh to represent the 4D temporal stages in the natural process. The constructed coarse dynamic mesh can be used to obtain the Neural Template as in Figure 4, which will be used in the second stage. We discuss the details of this process in Sec. 4.
@@ 4 Neural Template
4.1 Estimating Coarse 4D Geometry
@@ 5.1 Hybrid 4D Representation
We now introduce an implicit neural 4D representation for $\mathcal{G}$ that is both expressive and temporally consistent. The geometry of $\mathcal{G}$ is modeled as a time-evolving implicit level-set [61, 34, 38], represented by an implicit neural SDF. Given a spatial coordinate $\mathbf{x}$ and timestamp $t$, we compute a feature vector $f_{\text{4D}}(\mathbf{x},t)$, which is subsequently passed through two MLPs. These MLPs output SDF values and material parameters essential for neural rendering [60, 36] as detailed in Sec. 5.2.
The neural 4D feature vector $f_{\text{4D}}$ is computed via a hybrid 4D representation with K-Planes [13] representing low-frequency information and Neural Graphical Primitives (NGP) [37] representing high-frequency details.
@@
where $\psi_{t}$ is the NGP encoding for timestamp $t$, and $t_{i}$ and $t_{i+1}$ are two keyframes around the sampled timestamp. The full 4D feature $f_{\text{4D}}$ is computed by concatenating $f_{\text{low}}$ and $f_{\text{high}}$.
@@ 5.2 Physically-Based Rendering
We now describe the rendering procedure for the hybrid representation of $\mathcal{G}$ from Sec. 5.1.
Shading.
Following common conventions, we use the Disney Principled PBR material model [7] to perform the shading process. Specifically, the material parameters consist of diffuse color $k_{d}$, the visibility map $V(\mathbf{x})$, and $k_{\text{orm}}=(o,r,m)$ with roughness $r$, metallic parameter $m$, and $o$ left unused. The specular color $k_{s}$ is computed with $k_{s}=(1-m)\cdot 0.04+m\cdot k_{d}$. The outgoing radiance $L(\omega_{o})$ in direction $\omega_{o}$ can be represented with the rendering equation [22]:
@@
where $\mathbf{n}$ is the surface normal, $L_{i}(\mathbf{x},\omega_{i})$ is the incoming radiance from direction $\omega_{i}$, $f(\omega_{i},\omega_{o})$ is the BSDF computed from the material parameters.
Volume rendering.
After computing the shading value at each spatial coordinate $\mathbf{x}$ and timestamp $t$, we follow a standard differentiable rendering process [36, 60] to produce the rendered frame at time $t$. These frames are then concatenated to form a complete video $V_{G}$. More details on rendering can be found in the supplementary material.
@@ 6.3 Ablations and Validations
Figure 10: Ablation of Neural Template guidance.
Generation without Guidance. We show that using guidance during generation is necessary. Removing the Neural Template guidance in the pipeline leads to results in Figure 10, with visual artifacts of 3D inconsistency across views and the tiny motion across frames, suggesting the importance of the guidance.
@@ Acknowledgments.
The paper’s title is inspired by the artwork, Nacer y Morir de una Rosa, by Colombian artist Rosa Navarro, recently on exhibit at the Orange County Museum of Art in 2023. We thank Haian Jin, Zizhang Li, and Ruocheng Wang for insightful discussions and feedback on the paper. This work is in part supported by NSF RI #2211258, #2338203, ONR MURI N00014-22-1-2740, and Samsung. YZ is in part supported by the Stanford Interdisciplinary Graduate Fellowship.